OpenStudy (anonymous):

if ab=3 and ac=7 find bc

hartnn (hartnn):

in terms of a ?

OpenStudy (anonymous):

yes

hartnn (hartnn):

multiply them: (ab)(bc)= 3*7 = ?

OpenStudy (anonymous):

it has to be in simplest radical form

hartnn (hartnn):

(ab)(ac)= 3*7 =21 so a^2bc=21 bc= 21/a^2 or bc = 21a^(-2)